# Rich and Poor Texture Contrast: A Simple yet Effective Approach for AI-generated Image Detection

### 10 Dec 2023 v2

@@ Abstract
Recent generative models show impressive performance in generating photographic images. Humans can hardly distinguish such incredibly realistic-looking AI-generated images from real ones. AI-generated images may lead to ubiquitous disinformation dissemination. Therefore, it is of utmost urgency to develop a detector to identify AI-generated images. Most existing detectors suffer from sharp performance drops over unseen generative models. In this paper, we propose a novel AI-generated image detector capable of identifying fake images created by a wide range of generative models. Our approach leverages the inter-pixel correlation contrast between rich and poor texture regions within an image. Pixels in rich texture regions exhibit more significant fluctuations than those in poor texture regions. This discrepancy reflects that the entropy of rich texture regions is larger than that of poor ones. Consequently, synthesizing realistic rich texture regions proves to be more challenging for existing generative models. Based on this principle, we divide an image into multiple patches and reconstruct them into two images, comprising rich-texture and poor-texture patches respectively. Subsequently, we extract the inter-pixel correlation discrepancy feature between rich and poor texture regions. This feature serves as a universal fingerprint used for AI-generated image forensics across different generative models. In addition, we build a comprehensive AI-generated image detection benchmark, which includes 16 kinds of prevalent generative models, to evaluate the effectiveness of existing baselines and our approach. Our benchmark provides a leaderboard for follow-up studies. Extensive experimental results show that our approach outperforms state-of-the-art baselines by a significant margin. Our project: https://fdmas.github.io/AIGCDetect/
Figure 1: We conduct a comprehensive AI-generated image detection benchmark, including 16 kinds of prevalent generative models [19, 20, 3, 52, 6, 33, 21, 9, 31, 39, 15] and commercial APIs [50, 37] like Midjourney [30]. Multiple cutting-edge detectors [47, 11, 18, 28, 25, 46, 48, 32] are presented in the benchmark. We visualize the results with a radar chart. The concentric circles denote the detection accuracy. Our approach outperforms the state-of-the-art detector about 5% over average detection accuracy.
@@ 1 Introduction
In this paper, we design an AI-generated fake image detector capable of identifying various unseen source data, including GAN-based [19, 20, 3, 52, 52, 6, 33] and diffusion-based generative models [9, 31, 39, 15]. Specifically, we only train a classifier based on a training set generated by ProGAN [19], and it works effectively over various models like Stable Diffusion [39], BigGAN [3], Midjourney [30], etc. To accomplish this goal, we need to extract a universal fingerprint across various generative models rather than directly training a binary classifier from the spatial domain. We first recall the evolution of fake image composition. In the early stage, generative models can create high-quality facial images like CelebA set [27], but they falter in creating multiple categories images like ImageNet set [8]. The reason is that the distribution approximation based on random noise for facial images is much easier than that of multiple categories images. The entropy of the facial image distribution is much smaller than that of various categories images. It is difficult to approximate a distribution with high entropy from random noise. Modern state-of-the-art generative models can create realistic-looking fake images with arbitrary content. However, the poor texture regions of synthetic images still behave differently from rich texture regions. Generative models leave different artifacts between the poor and rich texture regions within synthetic images stemming from the entropy discrepancy (distribution approximation difficulty) of the two regions. We find that this characteristic of synthetic images widely exists across various cutting-edge generative models. Accordingly, we design an AI-generated fake image detector which works effectively across various generative models.
@@
  • Based on the universal fingerprint of synthetic images, we elaborate a novel AI-generated fake image detection approach capable of accurately identifying synthetic images generated by various generative models.
@@ 2 Related Work
Image generation has achieved a series of remarkable progress in recent years, which aims to generate photographic images from random noise [19, 20, 3, 52, 6, 33, 9] or text description [31, 15, 38, 35]. We denote the distribution of real images as $p(x)$. It is difficult to give a concrete formulaic expression of $p(x)$. In the image generation field, researchers aim to adopt $g_{\theta}(z)$ to approximate $p(x)$ as closely as possible, where $g_{\theta}(\cdot)$ and $z$ denote a learnable function (deep model) and random noise, respectively. For instance, famous GANs [19, 20, 3, 52, 6, 33] and their variants employ a generator to transfer random noise $z$ (sampled from an isotropic Gaussian distribution) to fake images, that is, mapping a Gaussian distribution into an image distribution. Then, they employ a discriminator, which aims to identify whether an image is real or fake, to measure the distance between the synthetic image distribution $g_{\theta}(z)$ and real image $p(x)$. The entire process of GANs consists of two steps, i.e., 1) measuring the distance between synthetic image distribution $g_{\theta}(z)$ and real image $p(x)$; 2) minimizing their distance by updating the generator $g_{\theta}(\cdot)$. In terms of prevalent diffusion models [39, 31, 9], they gradually introduce Gaussian noise to a real image, ultimately transforming it into an isotropic Gaussian noise. Subsequently, they learn to reverse the diffusion process in order to generate an image from a random noise sampled from an isotropic Gaussian distribution. In a nutshell, various image generation algorithms aim to approximate real image distribution $p(x)$ from an isotropic Gaussian distribution.
@@
AI-generated fake image detection aims to distinguish whether an image is real (generated by cameras or phones) or fake (generated by generative models). Prior studies [47, 11, 18, 28, 46, 48] treat fake image detection as a binary classification task. A deep model-based classifier (like ResNet-18) always performs well when the fake images of the test and training set are generated by the same generative model. However, its performance significantly degrades for unseen generative models. To tackle the generalization of the detector, many researchers have designed various approaches to explore a universal fingerprint across different generative models. Frank et al. [11] observe consistently anomalous behavior of fake images in the frequency domain and train the detector in the frequency domain. Durall et al. [10] further explain that the upsampling operation results in the frequency abnormality of fake images. Liu et al. [25] find that the noise pattern of an image in the frequency domain can be used as a fingerprint to improve the generalization of the detector. Liu et al. [28] focus on the texture of images and incorporate the Gram Matrix extraction into the typical ResNet structure to conduct fake image detection. Wang et al. [47] propose that easy data augmentation can significantly boost the generalization for the detector. More recently, Wang et al. [48] find that fake images generated by a diffusion model are more likely to be accurately reconstructed by a pre-trained diffusion model. Therefore, they adopt the reconstruction error of an image as the fingerprint to identify diffusion-generated fake images.
@@ 3.1 Motivation
As our aforementioned description, the goal of the image generation is to use an isotropic Gaussian distribution to approximate the real image distribution $p(x)$. The approximation difficulty depends on the entropy of the $p(x)$, which is related to the diversity of the real image $p(x)$. It is more difficult to approximate $p(x)$ with high diversity. This proposition is also corroborated by the development of the image generation field. In the early stage, generative models can create high-quality facial image, which is hardly distinguished from real facial images by humans, while their performance significantly drops when the training set is changed to a diversity set like ImageNet [8]. The real image distribution of ImageNet [8] (large entropy) is much more diverse than that of a facial set like CelebA [27].
Although existing state-of-the-art generative models can create realistic fake images with various content, they leave different artifacts in the rich or poor texture regions. Rich texture regions of an image behave more diversely than poor texture regions. As a result, it is more difficult for generative models to approximate rich texture regions of real images. An alternative crux of our approach is inter-pixel correlation analysis. Prior studies [11, 46, 25, 10] also show that the high frequency or noise of fake images exhibits large gaps from real images, which is conducive to improving the generalization of the detector. Therefore, we design a novel approach to extract inter-pixel correlation, regardless of the content of images. For real images, their inter-pixel correlations of different regions behave consistently. Based on the image device/tracking forensics [29], inter-pixel correlation (also is similar with PRNU noise pattern [5, 41]) is determined by its camera device (CMOS) and ISP (Image Signal Processing). Therefore, the inter-pixel correlations of real images between rich and poor texture regions are very close. Subsequently, we leverage the inter-pixel correlation contrast between rich and poor texture regions of an image as a fingerprint for generated images to identify AI-generated fake images. Fig. 2 illustrates our motivation.
@@ 3.2 Fingerprint Feature Extraction
Fig. 3 shows the framework of our approach, which consists of two steps, i.e., fingerprint feature extraction and detector training. Since our fingerprint feature is the inter-pixel correlation contrast between rich and poor texture regions, we aim to suppress the semantic information of images and extract inter-pixel correlation. Each input image is transferred into two parts, i.e., rich texture regions and poor texture regions. We first randomly crop the input image into multiple patches and sort patches based on their texture diversity. Then, we adopt these patches to reconstruct two images consisting of rich or poor texture patches, respectively. We name this process as Smash&Reconstruction, which is used to suppress the semantic information of the images. Smash&Reconstruction makes the detector distinguish real or fake images regardless of the semantic information of images. If the detector distinguishes real or fake images based on semantic information, its performance significantly drops for fake images whose semantic information is almost impeccable. The texture diversity can be measured by pixel fluctuation degree. Compared with rich texture regions, pixels located in the poor texture regions are prone to the same value. For a $M\times M$ size patch, we measure its texture diversity by summing the residual of four directions including diagonal, counter-diagonal, horizontal direction and vertical direction. Equation 1 gives the concrete expression,
@@
Patch images with rich texture are assigned to a large $l_{div}$. As Fig. 3 shows, the rich texture part consists of patches with high $l_{div}$, while the poor texture part consists of patches with low $l_{div}$. The semantic information of both two parts is broken due to Smash&Reconstruction.
@@ 4.1 Experimental Setup
DALL-E 2 [37] 256×256 2k ImageNet
@@
ProGAN 100.00 99.36 100.00 99.99 99.67 99.83 95.19 52.75 99.81 100.00
StyleGan 90.17 78.02 85.20 87.05 91.75 91.08 83.03 51.31 84.93 92.77
BigGAN 71.17 81.97 77.40 67.33 77.75 85.62 70.12 49.70 95.08 95.80
CycleGAN 87.62 78.77 87.00 86.07 84.10 86.94 74.19 49.58 98.33 70.17
StarGAN 94.60 94.62 97.00 95.05 99.92 99.27 95.47 46.72 95.75 99.97
GauGAN 81.42 80.57 77.00 69.35 75.39 78.46 67.79 51.23 99.47 71.58
Stylegan2 86.91 66.19 83.30 87.28 94.64 85.32 75.31 51.72 74.96 89.55
WFIR 91.65 50.75 66.80 86.80 70.85 55.70 58.05 53.30 86.90 85.80
ADM 60.39 63.42 49.00 58.61 84.73 67.15 75.78 98.25 66.87 82.17
Glide 58.07 54.13 57.20 54.50 80.52 66.11 71.75 92.42 62.46 83.79
Midjourney 51.39 45.87 52.20 50.02 65.55 65.35 58.01 89.45 56.13 90.12
SDv1.4 50.57 38.79 51.00 51.70 85.55 63.02 49.74 91.24 63.66 95.38
SDv1.5 50.53 39.21 51.40 52.16 85.67 63.67 49.83 91.63 63.49 95.30
VQDM 56.46 77.80 55.10 52.86 74.46 72.99 53.68 91.90 85.31 88.91
Wukong 51.03 40.30 51.70 50.76 82.06 59.55 54.46 90.90 70.93 91.07
DALLE2 50.45 34.70 52.80 49.25 88.75 65.45 66.48 92.45 50.75 96.60
Average 70.78 64.03 68.38 68.67 83.84 75.34 68.68 71.53 78.43 89.31
@@
Datasets In order to comprehensively evaluate our approach, we adopt 16 different generative models. The details of test datasets are shown in Table 1. The fake images generated by GAN-based generative models and whichfaceisreal (WFIR) [49] are contributed by CNNSpot [47]. Images of DALL-E 2 [37] are curated by ours. Other images are contributed by GenImage [53]. We use the label of the image as the prompt to generate fake images for some text2image generative models like DALL-E 2 and Midjourney. In terms of the training set, we adopt the training set proposed by CNNSpot [47], which is also widely used in previous studies [32, 46]. The training set contains 720k images made up of 360k real images from LSUN [51] and 360k fake images from ProGAN [19]. Since the detector does NOT know the generative model of fake images in the practical application, the generalization, which is the performance of the detector for unseen source data during the training phase, is an important criterion for the detector evaluation. As Table 1 shows, we adopt various generative models, including some advanced commercial APIs like Midjourney, to demonstrate the generalization of the proposed approach. Compared with DeepFake detection, which is dedicated to identifying face manipulation, we aim to identify AI-generated images with various semantic content. In other words, we focus on AI-synthesized image detection rather than AI-manipulated image detection. In terms of AI-generated facial images, We also evaluate detectors over facial images like whichfaceisreal (WFIR) [49] and StarGAN [6]. It is a challenging task that our training set does not contain facial images.

### 07 Mar 2024 v3

# PatchCraft: Exploring Texture Patch for Efficient AI-generated Image Detection

@@ Abstract
Recent generative models show impressive performance in generating photographic images. Humans can hardly distinguish such incredibly realistic-looking AI-generated images from real ones. AI-generated images may lead to ubiquitous disinformation dissemination. Therefore, it is of utmost urgency to develop a detector to identify AI-generated images. Most existing detectors suffer from sharp performance drops over unseen generative models. In this paper, we propose a novel AI-generated image detector capable of identifying fake images created by a wide range of generative models. We observe that the texture patches of images tend to reveal more traces left by generative models compared to the global semantic information of the images. A novel Smash&\&&Reconstruction preprocessing is proposed to erase the global semantic information and enhance texture patches. Furthermore, pixels in rich texture regions exhibit more significant fluctuations than those in poor texture regions. Synthesizing realistic rich texture regions proves to be more challenging for existing generative models. Based on this principle, we leverage the inter-pixel correlation contrast between rich and poor texture regions within an image to further boost the detection performance. In addition, we build a comprehensive AI-generated image detection benchmark, which includes 17 kinds of prevalent generative models, to evaluate the effectiveness of existing baselines and our approach. Our benchmark provides a leaderboard for follow-up studies. Extensive experimental results show that our approach outperforms state-of-the-art baselines by a significant margin. Our project: https://fdmas.github.io/AIGCDetect/
Figure 1: We conduct a comprehensive AI-generated image detection benchmark, including 16 kinds of prevalent generative models [19, 20, 3, 52, 6, 33, 21, 9, 31, 39, 15] and commercial APIs [50, 37] like Midjourney [30]. Multiple cutting-edge detectors [47, 11, 18, 28, 25, 46, 48, 32] are presented in the benchmark. We visualize the results with a radar chart. The concentric circles denote the detection accuracy. Our approach outperforms the state-of-the-art detector about 4% over average detection accuracy.
@@ 1 Introduction
In this paper, we design an AI-generated fake image detector capable of identifying various unseen source data, including GAN-based [19, 20, 3, 52, 52, 6, 33] and diffusion-based generative models [9, 31, 39, 15]. Specifically, we only train a classifier based on a training set generated by ProGAN [19], and it works effectively over various models like Stable Diffusion [39], BigGAN [3], Midjourney [30], etc. To accomplish this goal, we need to extract a universal fingerprint across various generative models rather than directly training a binary classifier from the spatial domain. Specifically, we identify AI-generated images based on the inter-pixel correlation. The inter-pixel correlation is closely related to the high-frequency component of an image. Previous studies [11, chandrasegaran2021closer, 10] demonstrate that various generative models leave many artifacts in high-frequency components of synthetic images rather than in low-frequency components related to the global semantic information of images. Therefore, we propose Smash&\&&Reconstruction to break the global semantic information (suppress low-frequency components) and magnify artifacts left by generative models. It is difficult to identify synthetic images from their semantic information for the state-of-the-art generative models. As a result, the performance of detectors based on semantic information sharply degrades for realistic-looking fake images. Smash&\&&Reconstruction makes the subsequent classifier distinguish real or synthetic images by the inter-pixel correlation extracted from texture patches instead of the global semantic information.
Next, We recall the evolution of fake image composition. In the early stage, generative models can create high-quality facial images like CelebA set [27], but they falter in creating multiple categories images like ImageNet set [8]. The reason is that the distribution approximation based on random noise for facial images is much easier than that of multiple categories images. The entropy of the facial image distribution is much smaller than that of various categories images. It is difficult to approximate a distribution with high entropy from random noise. Modern state-of-the-art generative models can create realistic-looking fake images with arbitrary content. However, the poor texture regions of synthetic images still behave differently from rich texture regions. Generative models leave different artifacts between the poor and rich texture regions within synthetic images stemming from the entropy discrepancy (distribution approximation difficulty) of the two regions. We find that this characteristic of synthetic images widely exists across various cutting-edge generative models. Accordingly, we design an AI-generated fake image detector which works effectively across various generative models.
@@
  • We propose Smash&\&&Reconstruction to break the semantic information of images and enhance the texture regions of images. Smash&\&&Reconstruction makes the detector focus on the inter-pixel correlation of images and significantly boosts the generalization of the detector.
@@ 2 Related Work
Image generation has achieved a series of remarkable progress in recent years, which aims to generate photographic images from random noise [19, 20, 3, 52, 6, 33, 9] or text description [31, 15, 38, 35]. We denote the distribution of real images as $p(x)$. It is difficult to give a concrete formulaic expression of $p(x)$. In the image generation field, researchers aim to adopt $g_{\theta}(z)$ to approximate $p(x)$ as closely as possible, where $g_{\theta}(\cdot)$ and $z$ denote a learnable function (deep model) and random noise, respectively. For instance, famous GANs [19, 20, 3, 52, 6, 33] and their variants employ a generator to transfer random noise $z$ (sampled from an isotropic Gaussian distribution) to fake images, that is, mapping a Gaussian distribution into an image distribution. Then, they employ a discriminator, which aims to identify whether an image is real or fake, to measure the distance between the synthetic image distribution $g_{\theta}(z)$ and real image $p(x)$. The entire process of GANs consists of two steps, i.e.,, 1) measuring the distance between synthetic image distribution $g_{\theta}(z)$ and real image $p(x)$; 2) minimizing their distance by updating the generator $g_{\theta}(\cdot)$. In terms of prevalent diffusion models [39, 31, 9], they gradually introduce Gaussian noise to a real image, ultimately transforming it into an isotropic Gaussian noise. Subsequently, they learn to reverse the diffusion process in order to generate an image from a random noise sampled from an isotropic Gaussian distribution. In a nutshell, various image generation algorithms aim to approximate real image distribution $p(x)$ from an isotropic Gaussian distribution.
@@
AI-generated fake image detection aims to distinguish whether an image is real (generated by cameras or phones) or fake (generated by generative models). Prior studies [47, 11, 18, 28, 46, 48] treat fake image detection as a binary classification task. A deep model-based classifier (like ResNet-50) always performs well when the fake images of the test and training set are generated by the same generative model. However, its performance significantly degrades for unseen generative models. To tackle the generalization of the detector, many researchers have designed various approaches to explore a universal fingerprint across different generative models. Frank et al. [11] observe consistently anomalous behavior of fake images in the frequency domain and train the detector in the frequency domain. Durall et al. [10] further explain that the upsampling operation results in the frequency abnormality of fake images. Liu et al. [25] find that the noise pattern of an image in the frequency domain can be used as a fingerprint to improve the generalization of the detector. Liu et al. [28] focus on the texture of images and incorporate the Gram Matrix extraction into the typical ResNet structure to conduct fake image detection. Wang et al. [47] propose that easy data augmentation can significantly boost the generalization for the detector. More recently, Wang et al. [48] find that fake images generated by a diffusion model are more likely to be accurately reconstructed by a pre-trained diffusion model. Therefore, they adopt the reconstruction error of an image as the fingerprint to identify diffusion-generated fake images.
@@ 3.1 Motivation
In this section, we aim to answer ”How to identify AI-generated images?”. A naive detection approach is training a common classifier like ResNet-50 to distinguish real and synthetic images directly. However, the performance of naive classifier always significantly degrades for unseen generative models. The reason is that the classifier largely relies on the global semantic information of images to make decisions, that is, the naive classifier can not catch the universal artifacts features of various generative models. Previous studies show that generative models leave artifacts in the high-frequency components due to the upsampling operation [10]. The abnormality in high-frequency components can result in the inter-pixel correlation discrepancy between synthetic and real images. Based on the image device/tracking forensics [29], inter-pixel correlation (also is similar with PRNU noise pattern [5, 41]) is determined by its camera device (CMOS) and ISP (Image Signal Processing). Although some cutting-edge generative models create impressive images from the semantic view, they still struggle to simulate inter-pixel correlation of real images. Therefore, we propose Smash&\&&Reconstruction to break the global semantic information and make the classifier focus on the inter-pixel correlations of images.
The goal of the image generation is to use an isotropic Gaussian distribution to approximate the real image distribution $p(x)$. The approximation difficulty depends on the entropy of the $p(x)$, which is related to the diversity of the real image $p(x)$. It is more difficult to approximate $p(x)$ with high diversity. This proposition is also corroborated by the development of the image generation field. In the early stage, generative models can create high-quality facial image, which is hardly distinguished from real facial images by humans, while their performance significantly drops when the training set is changed to a diversity set like ImageNet [8]. The real image distribution of ImageNet [8] (large entropy) is much more diverse than that of a facial set like CelebA [27].
Although existing state-of-the-art generative models can create realistic fake images with various content, they leave different artifacts in the rich or poor texture regions. Rich texture regions of an image behave more diversely than poor texture regions. As a result, it is more difficult for generative models to approximate rich texture regions of real images. For a real image, the inter-pixel correlations, which are determined by its camera device (CMOS) and ISP (Image Signal Processing), between rich and poor texture regions are very close. Therefore, we leverage the inter-pixel correlation contrast between rich and poor texture regions of an image as a fingerprint to identify AI-generated fake images. Fig. 2 illustrates our motivation of the inter-pixel correlation contrast.
@@ 3.2 Fingerprint Feature Extraction
Fig. 3 shows the framework of our approach, which consists of two steps, i.e.,, fingerprint feature extraction and detector training. Since our fingerprint feature is the inter-pixel correlation contrast between rich and poor texture regions, we aim to suppress the semantic information of images and extract inter-pixel correlation. Each input image is transferred into two parts, i.e., rich texture regions and poor texture regions. We first randomly crop the input image into multiple patches and sort patches based on their texture diversity. Then, we adopt these patches to reconstruct two images consisting of rich or poor texture patches, respectively. We name this process as Smash&\&&Reconstruction, which is used to suppress the semantic information of the images. Smash&\&&Reconstruction makes the detector distinguish real or fake images regardless of the semantic information of images. If the detector distinguishes real or fake images based on semantic information, its performance significantly drops for fake images whose semantic information is almost impeccable. The texture diversity can be measured by pixel fluctuation degree. Compared with rich texture regions, pixels located in the poor texture regions are prone to the same value. For a $M\times M$ size patch, we measure its texture diversity by summing the residual of four directions including diagonal, counter-diagonal, horizontal direction and vertical direction. Equation 1 gives the concrete expression,
@@
Patch images with rich texture are assigned to a large $l_{div}$. As Fig. 3 shows, the rich texture part consists of patches with high $l_{div}$, while the poor texture part consists of patches with low $l_{div}$. The semantic information of both two parts is broken due to Smash&\&&Reconstruction.
@@ 4.1 Experimental Setup
DALL-E 2 [37] 256×256 2.0k ImageNet
SDXL [podell2023sdxl] 1024×1024 4.0k COCO
@@
ProGAN 100.00 99.36 100.00 99.99 99.95 99.83 95.19 52.75 99.81 100.00
StyleGan 90.17 78.02 85.20 87.05 92.64 91.08 83.03 51.31 84.93 92.77
BigGAN 71.17 81.97 77.40 67.33 88.43 85.62 70.12 49.70 95.08 95.80
CycleGAN 87.62 78.77 87.00 86.07 79.07 86.94 74.19 49.58 98.33 70.17
StarGAN 94.60 94.62 97.00 95.05 100.00 99.27 95.47 46.72 95.75 99.97
GauGAN 81.42 80.57 77.00 69.35 79.17 78.46 67.79 51.23 99.47 71.58
Stylegan2 86.91 66.19 83.30 87.28 93.82 85.32 75.31 51.72 74.96 89.55
whichfaceisreal 91.65 50.75 66.80 86.80 50.00 55.70 58.05 53.30 86.90 85.80
ADM 60.39 63.42 49.00 58.61 83.91 67.15 75.78 98.25 66.87 82.17
Glide 58.07 54.13 57.20 54.50 83.50 66.11 71.75 92.42 62.46 83.79
Midjourney 51.39 45.87 52.20 50.02 69.55 65.35 58.01 89.45 56.13 90.12
SDv1.4 50.57 38.79 51.00 51.70 89.33 63.02 49.74 91.24 63.66 95.38
SDv1.5 50.53 39.21 51.40 52.16 88.81 63.67 49.83 91.63 63.49 95.30
VQDM 56.46 77.80 55.10 52.86 85.03 72.99 53.68 91.90 85.31 88.91
wukong 51.03 40.30 51.70 50.76 86.39 59.55 54.46 90.90 70.93 91.07
DALLE2 50.45 34.70 52.80 49.25 92.45 65.45 66.48 92.45 50.75 96.60
SDXL 53.03 51.23 55.60 64.53 87.75 71.30 55.35 91.28 50.73 98.43
Average 69.73 63.28 67.63 68.43 85.28 75.11 67.90 72.70 76.80 89.85
@@
Datasets. In order to comprehensively evaluate our approach, we adopt 17 different generative models. The details of test datasets are shown in Table 1. The fake images generated by GAN-based generative models and whichfaceisreal (WFIR) [49] are contributed by CNNSpot [47]. Images of DALL-E 2 [37] and SDXL [podell2023sdxl] are curated by ours. Other images are contributed by GenImage [53]. We use the label of the image as the prompt to generate fake images for some text2image generative models like DALL-E 2 and Midjourney. In terms of the training set, we adopt the training set proposed by CNNSpot [47], which is also widely used in previous studies [32, 46]. The training set contains 720k images made up of 360k real images from LSUN [51] and 360k fake images from ProGAN [19]. Since the detector does NOT know the generative model of fake images in the practical application, the generalization, which is the performance of the detector for unseen source data during the training phase, is an important criterion for the detector evaluation. As Table 1 shows, we adopt various generative models, including some advanced commercial APIs like Midjourney, to demonstrate the generalization of the proposed approach. Compared with DeepFake detection, which is dedicated to identifying face manipulation, we aim to identify AI-generated images with various semantic content. In other words, we focus on AI-synthesized image detection rather than AI-manipulated image detection. In terms of AI-generated facial images, We also evaluate detectors over facial images like whichfaceisreal (WFIR) [49] and StarGAN [6]. It is a challenging task that our training set does not contain facial images.